OpenStudy (anonymous):

Create your own third degree polynomial that when divided by x + 2 has a remainder of –4

OpenStudy (valpey):

Poly(ax^3+bx^2+cx+d)+4=(x+2)*Poly(ex^2+fx+g)

OpenStudy (valpey):

If you want to keep it simple let e =1 and f and g = 0 then: Poly(ax^3+bx^2+cx+d)+4=(x+2)*x^2=x^3+2x^2 That would imply a=1 and b=2 and c=0 Then d = -4

OpenStudy (valpey):

So \[x^3+2x^2-4\] is one such polynomial.